### v1

@@ 3 Quasi-orthogonality
(10) $\operatorname{dist}_{a}(x_{ij},y_{ij})\leqslant\left(1+\frac{2}{\theta^{d_{i}-3}}\left(\left(1+\theta\right)^{d_{i}-1}-\theta^{d_{i}-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}}\triangleq\tilde{\varepsilon}.$
@@
Set $\{\beta_{rt}^{(ij)}\}$ such that $y_{ij}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$ and we have that
@@
$B_{1}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(1d_{1})}&\cdots&\beta_{11}^{(q1)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{1d_{1}}^{(11)}&\cdots&\beta_{1d_{1}}^{(1d_{1})}&\cdots&\beta_{1d_{1}}^{(q1)}&\cdots&\beta_{1d_{1}}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{p1}^{(11)}&\cdots&\beta_{p1}^{(1d_{1})}&\cdots&\beta_{p1}^{(q1)}&\cdots&\beta_{p1}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{pd_{p}}^{(11)}&\cdots&\beta_{pd_{p}}^{(1d_{1})}&\cdots&\beta_{pd_{p}}^{(q1)}&\cdots&\beta_{pd_{p}}^{(qd_{q})}\end{pmatrix}.$
We affirm that matrix
$B_{2}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots\\ \beta_{qd_{q}}^{(11)}&\cdots&\beta_{qd_{q}}^{(qd_{q})}\end{pmatrix},$
is strictly diagonally dominant provided $\tilde{\varepsilon}<\frac{1}{\sqrt{N}}$. In fact, we obtain from Eq. 10 that
$\tilde{\varepsilon}\geqslant\operatorname{dist}_{a}(x_{ij},y_{ij})=\operatorname{dist}_{a}(y_{ij},x_{ij})=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}-\beta_{ij}^{(ij)}x_{ij}\right\|_{a}}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}\right\|_{a}}.$
@@
and $\operatorname{rank}(B_{2})=N=\operatorname{rank}(B_{1})$, which completes the proof.
Back to Eq. 3, we turn to estimate the distance between the orthonormal approximate eigenfunctions and the orthonormal exact eigenfunctions.
Theorem 3.7.
If $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, then there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{kl})=\delta_{ik}\delta_{jl}$ such that
$\operatorname{dist}_{a}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C$ is a constant that is independent of $V^{h}$.
Proof 3.8.
For the approximate eigenpairs $\left\{(\lambda^{h}_{ij},u^{h}_{ij})\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}$, we obtain from Proposition 2.4 that there exists $\hat{u}_{ij}\in M(\lambda_{i})$ such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}\leqslant L\operatorname{dist}_{a}(M(\lambda_{i}),V^{h})\leqslant L\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1,\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $L$ is a constant that is independent of $V^{h}$, which yields that
$\left\|\frac{u^{h}_{ij}}{\|u^{h}_{ij}\|_{a}}-\frac{\hat{u}_{ij}}{\|\hat{u}_{ij}\|_{a}}\right\|_{a}=\frac{1}{\sqrt{\lambda^{h}_{ij}}}\left\|u^{h}_{ij}-\hat{u}_{ij}+\left(1-\frac{\|u^{h}_{ij}\|_{a}}{\|\hat{u}_{ij}\|_{a}}\right)\hat{u}_{ij}\right\|_{a}$
$\leqslant$ $\frac{1}{\sqrt{\lambda}_{i}}\left(\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}+\left|\|\hat{u}_{ij}\|_{a}-\|u^{h}_{ij}\|_{a}\right|\right)\leqslant\frac{2}{\sqrt{\lambda}_{i}}L\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right).$
Then it follows from Corollary 3.3 that there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{ik})=\delta_{jk}$ such that
$\operatorname{dist}_{a}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad j=1,\ldots,d_{i},$
where $C$ is a constant that is independent of $V^{h}$. By traversing $i=1,\ldots,q$, the proof is completed.
We see that Theorem 3.7 tells that, for the finite dimensional approximation of an eigenvalue problem, there exists a set of orthogonal eigenfunctions whose distance to the orthogonal approximate eigenfunctions is controlled by the distance of subspaces.
In next section, we will present the approximation between iterative solutions $\left\{\left(\lambda_{ij}^{(n)},u_{ij}^{(n)}\right)\right\}$ produced by the ParO approach and solutions of the discrete problem Eq. 3 with the application of Theorem 3.7. Then we obtain the approximation errors between iterative solutions and solutions of Eq. 1 from Proposition 2.3, Theorem 3.7 and the triangle inequalities
$\operatorname{dist}_{a}(u_{ij}^{o},u_{ij}^{(n)})$ $\leqslant\operatorname{dist}_{a}(u_{ij}^{o},u^{h,o}_{ij})+\operatorname{dist}_{a}(u^{h,o}_{ij},u_{ij}^{(n)}),$
4 Numerical Analysis
With the quasi-orthogonality, in this section, we carry out the numerical analysis of the ParO approach for clustered eigenvalue problems.
4.1 Algorithm framework
We first recall the framework of the ParO approach for the first $N$ clustered eigenvalues and their corresponding eigenfunctions of Eq. 1, which is stated as Algorithm 1. We mention that Algorithm 1 is indeed a modified version of Algorithm 1.1 in [9].
Algorithm 1 A framework for the ParO approach
  1. 1.
    Given a finite dimensional subspace $V^{h}$ and initial data $\left(\lambda_{k}^{(0)},u_{k}^{(0)}\right)\in\mathbb{R}\times V^{h}$ with $b\left(u_{i}^{(0)},u_{j}^{(0)}\right)=\delta_{ij}(i,j=1,2,\ldots,N)$, let $n=0$.
  2. 2.
    For $k=1,2,\ldots,N$, update each orbital $u_{k}^{(n)}$ in parallel and obtain $u_{k}^{(n+1/2)}$.
  3. 3.
    Construct $U_{n+1}=\operatorname{span}\left\{u_{1}^{(n+1/2)},u_{2}^{(n+1/2)},\ldots,u_{N}^{(n+1/2)}\right\}$.
  4. 4.
    If necessary, find a new basis of $U_{n+1}$ by some procedure and obtain eigenpairs $\left(\lambda_{k}^{(n+1)},u_{k}^{(n+1)}\right)$ by some way, or let $u_{k}^{(n+1)}=u_{k}^{(n+1/2)}$ for $k=1,\ldots,N$.
  5. 5.
    If not converge, let $n=n+1$ and go to $2$.
  • solving a coarse eigenvalue problem as follows: given a finite dimensional subspace $V^{H}$ of $H$ with $\operatorname{dim}(V^{H})>N$, find $(\lambda^{H},u^{H})\in\mathbb{R}\times V^{H}$ satisfying
    (11) $a\left(u^{H},v\right)=\lambda^{H}b\left(u^{H},v\right)\quad\forall v\in V^{H},$
    to obtain eigenpairs $(\lambda_{k}^{H},u_{k}^{H})$ satisfying $b\left(u_{i}^{H},u_{j}^{H}\right)=\delta_{ij}$ for $i,j=1,2,\ldots,N$ and set $(\lambda_{k}^{(0)},u_{k}^{(0)})=(\lambda_{k}^{H},u_{k}^{H})$ for $k=1,\ldots,N$;
  • neural networks based guesses, which can be obtained from the subspace method based on neural networks [30].
Note that eigenvalue problems are usually resulting from physics. We are able to apply the initial values from physical observation or data. For instance, as mentioned in [9], in electronic structure calculations, we may choose initial data from Gaussian-type orbital, Slater-type orbital and atomic orbital based guesses, and so on.
Since we look for clustered eigenvalues and their corresponding eigenfunctions, we shall consider the approximation of each eigenspace as mentioned in Introduction. We understand that it is not trivial to obtain the multiplicity $d_{i}$ of each eigenvalue $\lambda_{i}$. An effective way to approximate the multiplicities is to cluster the initial guesses $\lambda_{1}^{(0)}\leqslant\lambda_{2}^{(0)}\leqslant\cdots\leqslant\lambda_{N}^{(0)}$. By clustering methods such as Bayesian Information Criterion and Silhouette Method (see e.g., [26, 28]), we can get $q^{\prime}$ clusters with $d_{i}^{\prime}$ eigenpairs in the $i-$th cluster $(i=1,\ldots,q^{\prime})$, that is,
In this papaer, we assume that $q^{\prime}=q$ and $d_{i}^{\prime}=d_{i}$ for $i=1,\ldots,q^{\prime}$. Indeed, with a sufficient a priori information of the eigenvalue problem, such an assumption will be likely to hold.
Set
then $U_{n}=\sum_{i=1}^{q}U_{n}^{(i)}$.
To update each orbital in step $2$ and step $3$ of Algorithm 1, as pointed in [9], we can apply the shifted-inverse approach, Chebyshev filtering and so on. Define $\mathcal{F}_{n}:U_{n}\rightarrow U_{n+1}$ as follows
$\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:u_{ij}^{(n)}=u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n)}\mapsto u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n+1/2)}\triangleq u_{ij}^{(n+1/2)},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
Set
By step $3$, we have $U_{n+1}=\sum_{i=1}^{q}U_{n+1/2}^{(i)}$.
In step $4$ of Algorithm 1, the procedure to update the basis of $U_{n+1}$ can be using the (Gram-Schmidt) orthogonalization or solving a small scale eigenvalue problem as follows: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}$ satisfying
to obtain eigenpairs $(\lambda_{ij}^{(n+1)},u_{ij}^{(n+1)})$ with $b\left(u_{ij}^{(n+1)},u_{kl}^{(n+1)}\right)=\delta_{ik}\delta_{jl}$ for $(1,1)\leqslant(i,j),(k.l)\leqslant(q,d_{q})$.
The following theorem shows the approximation errors of eigenpairs when solving a small scale eigenvalue problem is carried out under the assumption that orbitals are approximated well.
Theorem 4.1.
If $\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)})\ll 1(i=1,\ldots,q)$, then after solving a small scale eigenvalue problem in $U_{n_{0}+1}=\sum_{i=1}^{q}U_{n_{0}+1/2}^{(i)}$, there exists an orthonormal basis $\{u^{h,o}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o}_{ij},u^{h,o}_{kl})=\delta_{ik}\delta_{jl}$ such that for $(i,j)\leqslant(q,d_{q})$
(15) $|\lambda_{ij}^{(n_{0}+1)}-\lambda^{h}_{ij}|\leqslant$ $L_{1}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}^{2}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
(16) $\operatorname{dist}_{a}(u^{h,o}_{ij},u_{ij}^{(n_{0}+1)})\leqslant$ $L_{2}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
where $L_{1}$ and $L_{2}$ are constants that are independent of $U_{n_{0}+1/2}$.
Proof 4.2.
@@ 4.1 Algorithm framework
We obtain from Proposition 3.5 that
$U_{n_{0}+1}=\bigoplus_{i=1}^{q}U_{n_{0}+1/2}^{(i)}.$

### v2

@@ 4.2 Shifted-inverse based ParO algorithm
$\operatorname{dist}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}:=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{\delta_{0}}{g-\delta_{0}}.$
Theorem 4.3 shows that convergence of Algorithm 2 does not require sufficiently accurate initial guesses. Eq. 10 ensures that the shift is closer to the eigenvalue being approximated. Indeed, Eq. 10 can be satisfied under the assumption that the finite dimensional discretization Eq. 3 approximates Eq. 1 not “too badly”. Eq. 11 guarantees that the dimension of the approximated subspace is preserved. When $N=1$, Eq. 30 may be reviewed as the classical shift-inverse convergence result.
Shifted-inverse based ParO algorithm
Algorithm 3 Shifted-inverse based ParO algorithm
  1. 1.
    Given a finite dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 6, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$. Set $\bar{\lambda}_{i}^{(0)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
  3. 3.
    Solve an eigenvalue problem: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}=\mathbb{R}\times\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)}\right\}$ satisfying
  4. 4.
    If $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left|\lambda_{ij}^{(n+1)}-\lambda_{ij}^{(n)}\right|>tol$, set $\bar{\lambda}_{i}^{(n+1)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(n+1)}\right\}_{j=1}^{d_{i}}\right)$, $n=n+1$ and go to Step $2$.
As mentioned above, we assume that the shifts are always not equal to any eigenvalue of Eq. 3 in the calculation process. If there are cases where some eigenfunctions are very well approximated, while other eigenfunctions have not yet converged, then we continue the iterative process on the non-converged eigenfunctions while keeping the well approximated eigenfunctions unchanged.
The following theorem tells the convergence of Algorithm 3. The proof is given in Section A.6
Theorem 4.4.
(14) $\operatorname{dist}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)$ $\leqslant\varepsilon_{0},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
$\operatorname{dist}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)$ $\leqslant\varepsilon_{n+1}:=\frac{\tilde{C}_{**}\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
$|\lambda^{h}_{ij}-\lambda_{ij}^{(n+1)}|$ $\leqslant\zeta_{n+1}:=\frac{\lambda^{h}_{q+1,1}}{\tilde{C}_{**}^{2}}\varepsilon_{n+1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $D:=\max_{1\leqslant i\leqslant q}d_{i}$ and the constant $\tilde{C}_{**}$ is defined in Theorem 4.1 and is is independent of $U_{n}(n=0,1,2,\ldots)$, and
(16) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{\tilde{C}_{**}\sqrt{DN}\gamma}{g-\gamma}.$
Note that $\gamma\ll 1$ when $\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, which together with Eq. 16 implies that Algorithm 3 converges faster when the finite dimensional discretization Eq. 3 approximates Eq. 1 better.
(17) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}^{3}}=\frac{\sqrt{DN}\lambda^{h}_{q+1,1}}{g\tilde{C}_{**}},$
which implies that it is a cubic convergence result. Note that the above cubic convergence stems from the convergence of the shifts to some eigenvalues of (2.3), which is exactly what $\gamma=0$ means. The constant $\tilde{C}_{**}$ in Eq. 17 comes from the application of Theorem 4.1. Indeed, we can choose a larger $\tilde{C}_{**}$ since Eq. 17 implies that Algorithm 3 converges faster with larger $\tilde{C}_{**}$. However, when $\tilde{C}_{**}$ is larger, more exact initial values are required. For example, if we expect $\varepsilon_{n}$ to decrease towards 00 as $n\rightarrow\infty$, a necessary condition
$\varepsilon_{0}\geqslant\varepsilon_{1}=\frac{\tilde{C}_{**}\sqrt{DN}\zeta_{0}\varepsilon_{0}}{\sqrt{(g-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\zeta_{0}^{2}\varepsilon^{2}_{0}}},$
implies that $\varepsilon_{0}$ and $\zeta_{0}$ are required smaller with larger $\tilde{C}_{**}$. Hence, from Theorem 3.7 and Eq. 16, we obtain that Algorithm 3 applied to a discrete eigenvalue problem converges in cubic rate and goes faster with more exact initial values. The classical result in the discrete case $(N=1)$ is stated as
Modified shifted-inverse based ParO algorithm
Algorithm 4 Modified Shifted-Inverse Based ParO Algorithm
  1. 1.
    Given a finite dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 6, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$. Set $\bar{\lambda}_{i}^{(0)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
  3. 3.
    Solve an eigenvalue problem: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times\tilde{U}_{n+1}=\mathbb{R}\times\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)},e_{11}^{(n+1/2)},\ldots,e_{qd_{q}}^{(n+1/2)}\right\}$ satisfying
  4. 4.
    If $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left|\lambda_{ij}^{(n+1)}-\lambda_{ij}^{(n)}\right|>tol$, set $\bar{\lambda}_{i}^{(n+1)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(n+1)}\right\}_{j=1}^{d_{i}}\right)$, $n=n+1$ and go to Step $2$.
$\tilde{U}_{n+1}=U_{n+1}\cup U_{n},\quad\operatorname{dist}(M_{h}(\lambda_{i}),\tilde{U}_{n+1})\leqslant\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1}).$
[…]
5 Concluding Remarks
In this paper, we have provided the numerical analysis of ParO for linear eigenvalue problems based on the investigation of a quasi-orthogonality. Within the framework of ParO, we have demonstrated the convergence of the associated practical algorithms. We point out that numerical experiments in [10, 12, 24] show that ParO is very efficient for electronic structure calculations. Due to the space limitation, we shall address the numerical analysis of the approach for the Kohn-Sham equation in a separate article. It is also our ongoing work to carry out the numerical analysis for the ParO based optimization approach proposed in [12].
Acknowledgments
The authors would like to thank Prof. Zhaojun Bai, and the anonymous referees for their constructive suggestions for the enhancement of the conclusions in Section 3 and the improvement of the presentation.
Appendix A Detailed Proofs
A.1 Proof of Proposition 3.2
Proof A.1.
For convenience, we first introduce the following notation. For $\{x_{i}\}_{i=1}^{n},\{y_{i}\}_{i=1}^{n}\subset H$, we set $\mathcal{X}=(x_{1},x_{2},\ldots,x_{n}),$ $\mathcal{Y}=(y_{1},y_{2},\ldots,y_{n})\in H^{n}$ and define
$\|\mathcal{X}\|=\sqrt{\sum_{i=1}^{n}\|x_{i}\|^{2}},\quad\text{and}\quad\mathcal{X}^{T}\mathcal{Y}=\begin{pmatrix}a(x_{1},y_{1})&\cdots&a(x_{1},y_{n})\\ \vdots&\ddots&\vdots\\ a(x_{n},y_{1})&\cdots&a(x_{n},y_{n})\end{pmatrix}.$
Let $\mathcal{V}=(v_{1},v_{2},\ldots,v_{n})$, since $\{v_{j}\}_{j=1}^{n}\subset H$ is $\delta$-quasi-orthogonal, there exists $\mathcal{U}=(u_{1},u_{2},\ldots,u_{n})$ satisfying that
(20) $\mathcal{U}^{T}\mathcal{U}=I,\quad\|\mathcal{U}-\mathcal{V}\|\leqslant\sqrt{\sum_{i=1}^{n}\|u_{i}-v_{i}\|^{2}}\leqslant\sqrt{n}\delta.$
Let $\operatorname{span}(\mathcal{V})=\operatorname{span}\{v_{1},\ldots,v_{n}\}$. If $\mathcal{W}\in H^{n}$ with $\mathcal{W}^{T}\mathcal{W}=I$ and $\operatorname{span}(\mathcal{W})=\operatorname{span}(\mathcal{V})$, then there exists the polar decomposition of $\mathcal{W}^{T}\mathcal{V}$ as follows,
(21) $\mathcal{W}^{T}\mathcal{V}=QP,$
where $Q\in\mathbb{R}^{n\times n}$ is an orthogonal matrix and $P\in\mathbb{R}^{n\times n}$ is positive definite. Denote $\tilde{\mathcal{V}}=\mathcal{W}Q$, and we see that $\tilde{\mathcal{V}}^{T}\tilde{\mathcal{V}}=Q^{T}\mathcal{W}^{T}\mathcal{W}Q=I$ and
(22) $\mathcal{V}=\mathcal{W}\mathcal{W}^{T}\mathcal{V}=\mathcal{W}QP=\tilde{\mathcal{V}}P.$
The decomposition Eq. 22 is unique since Eq. 21 implies
$P=\sqrt{(\mathcal{W}^{T}\mathcal{V})^{T}\mathcal{W}^{T}\mathcal{V}}=\sqrt{\mathcal{V}^{T}\mathcal{W}^{T}\mathcal{W}\mathcal{V}}=\sqrt{\mathcal{V}^{T}\mathcal{V}}.$
Next, we show that the maximum singular value $\sigma_{\text{max}}(\mathcal{U}^{T}\tilde{\mathcal{V}})\leqslant 1$. Consider $\mathcal{Z}\in H^{m}(m\geqslant n)$ with $\mathcal{Z}^{T}\mathcal{Z}=I$ and $\operatorname{span}(\mathcal{Z})=\operatorname{span}(\mathcal{U})+\operatorname{span}(\tilde{\mathcal{V}})$. We observe that there exist $Q_{1},Q_{2}\in\mathbb{R}^{m\times n}$ with $Q_{1}^{T}Q_{1}=Q_{2}^{T}Q_{2}=I$ satisfying
$\mathcal{U}=\mathcal{Z}Q_{1},\quad\tilde{\mathcal{V}}=\mathcal{Z}Q_{2},$
and
(23) $\sigma_{\text{max}}(\mathcal{U}^{T}\tilde{\mathcal{V}})=\sigma_{\text{max}}((\mathcal{Z}Q_{1})^{T}\mathcal{Z}Q_{2})=\sigma_{\text{max}}(Q_{1}^{T}Q_{2})\leqslant\sigma_{\text{max}}(Q_{1})\sigma_{\text{max}}(Q_{2})\leqslant 1.$
It follows from Eq. 22 and Eq. 23 that
$\|\mathcal{U}-\mathcal{V}\|^{2}=$ $\|\mathcal{U}\|^{2}+\|\mathcal{V}\|^{2}-2\operatorname{tr}(\mathcal{U}^{T}\mathcal{V})\geqslant\|\tilde{\mathcal{V}}\|^{2}+\|\mathcal{V}\|^{2}-2\sum_{i=1}^{n}\sigma_{i}(\mathcal{U}^{T}\mathcal{V})$
$=$ $\|\tilde{\mathcal{V}}\|^{2}+\|\mathcal{V}\|^{2}-2\sum_{i=1}^{n}\sigma_{i}(\mathcal{U}^{T}\tilde{\mathcal{V}}\tilde{\mathcal{V}}^{T}\mathcal{V})$
$\geqslant$ $\|\tilde{\mathcal{V}}\|^{2}+\|\mathcal{V}\|^{2}-2\sum_{i=1}^{n}\sigma_{\text{max}}(\mathcal{U}^{T}\tilde{\mathcal{V}})\sigma_{i}(P)$
$\geqslant$ $\|\tilde{\mathcal{V}}\|^{2}+\|\mathcal{V}\|^{2}-2\operatorname{tr}(\tilde{\mathcal{V}}^{T}\mathcal{V})=\|\tilde{\mathcal{V}}-\mathcal{V}\|^{2}.$
Hence, we have
$\|\tilde{\mathcal{V}}-\mathcal{V}\|\leqslant\|\mathcal{U}-\mathcal{V}\|,$
which together with Eq. 20 yields that for $\tilde{\mathcal{V}}=(\tilde{v}_{1},\tilde{v}_{2},\ldots,\tilde{v}_{n})$,
$\|\tilde{v}_{j}-v_{j}\|\leqslant\|\tilde{\mathcal{V}}-\mathcal{V}\|\leqslant\|\mathcal{U}-\mathcal{V}\|\leqslant\sqrt{n}\delta,\quad j=1,2,\ldots,n.$
A.2 Proof of Proposition 3.6
Proof A.2.
Let $\{x_{ij}\}_{j=1}^{d_{i}}$ be an orthonormal basis of $X_{i}(i=1,\ldots,p)$ with $a(x_{ij},x_{kl})=\delta_{ik}\delta_{jl}$. We obtain from Proposition 3.4 that there exists an orthonormal basis $\{y_{ij}\}_{j=1}^{d_{i}}$ of $Y_{i}(i=1,\ldots,q)$ satisfying that for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$,
(24) $\operatorname{dist}(x_{ij},y_{ij})\leqslant\max_{i=1,\ldots,q}(1+\sqrt{d_{i}})\sqrt{2-2\sqrt{1-\varepsilon^{2}}}:=\tilde{\varepsilon}.$
@@ Appendix A Detailed Proofs
Obviously, $\varepsilon<\min_{1\leqslant i\leqslant q}\sqrt{\frac{4(1+\sqrt{d_{i}})^{2}N-1}{4(1+\sqrt{d_{i}})^{4}N^{2}}}$ implies that $\tilde{\varepsilon}<\frac{1}{\sqrt{N}}$. Set $\{\beta_{rt}^{(ij)}\}$ such that $y_{ij}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$ and we have that
@@
$B_{1}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(1d_{1})}&\cdots&\beta_{11}^{(q1)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{1d_{1}}^{(11)}&\cdots&\beta_{1d_{1}}^{(1d_{1})}&\cdots&\beta_{1d_{1}}^{(q1)}&\cdots&\beta_{1d_{1}}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{p1}^{(11)}&\cdots&\beta_{p1}^{(1d_{1})}&\cdots&\beta_{p1}^{(q1)}&\cdots&\beta_{p1}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{pd_{p}}^{(11)}&\cdots&\beta_{pd_{p}}^{(1d_{1})}&\cdots&\beta_{pd_{p}}^{(q1)}&\cdots&\beta_{pd_{p}}^{(qd_{q})}\end{pmatrix}:=\begin{pmatrix}B_{2}\\ \star\end{pmatrix},$
with
$B_{2}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots\\ \beta_{qd_{q}}^{(11)}&\cdots&\beta_{qd_{q}}^{(qd_{q})}\end{pmatrix}.$
We assert that $B_{2}$ is strictly diagonally dominant. In fact, we obtain from Eq. 24 that
$\tilde{\varepsilon}\geqslant\operatorname{dist}(x_{ij},y_{ij})=\operatorname{dist}(y_{ij},x_{ij})=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}-\beta_{ij}^{(ij)}x_{ij}\right\|}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}\right\|}.$
@@
and $\operatorname{rank}(B_{2})=N$. Therefore, $\operatorname{rank}(B_{1})=N$, which completes the proof.
A.3 Proof of Theorem 4.1
Proof A.3.
@@
Since $\operatorname{dist}(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)})\ll 1(i=1,\ldots,q)$, we obtain from Proposition 3.6 that
$U_{n+1}=\bigoplus_{i=1}^{q}U_{n+1/2}^{(i)}.$